# Exponential actions defined by vector configurations, Gale duality, and moment-angle manifolds

### v1

@@ 8.3. From quadrics to polyhedra
Furthermore, the data $\{\mathcal{K},\mathrm{A}\}$ defines the normal fan of a generic polyhedron $P$ given by (8.1), where $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ is the Gale dual configuration in $W^{*}$ and the numbers $b_{i}$ are defined by the expression $\delta=\sum_{i=1}^{m}b_{i}\gamma_{i}$.
@@
By definition, $\mathcal{K}$ is a pure simplicial complex of dimension $m-k-1$. Condition (b) implies that if $\delta\in\mathop{\mathrm{cone}}\Gamma_{J}$, then $\delta\in\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{J}$, $\mathop{\mathrm{cone}}\Gamma_{J}$ is a full-dimensional cone (of dimension $k$) and its relative interior is the interior. Indeed, suppose $\delta\notin\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{J}$, then $\delta$ lies in a face of $\mathop{\mathrm{cone}}\Gamma_{J}$, which is a cone of dimension $<k$. By the Carathéodory Theorem [37, Proposition 1.15], this cone is generated by subset of $\Gamma$ of cardinality $<k$, in contradiction with (b).
@@
The decomposition of $\mathop{\mathrm{cone}}\Gamma$ into cones (or chambers) $C(\delta)$ is called the GKZ decomposition (after Gelfand, Kapranov and Zelevinsky), and $\Sigma_{\Gamma}$ is called the GKZ fan, or the secondary fan of $\Gamma$, see [19, Chapter 7] and also [29], [1, §2.2.2]. It encodes generalised normal fans with cone generators chosen among the Gale dual configuration $\mathrm{A}$. The relative interiors of maximal cones of $\Sigma_{\Gamma}$ correspond to the normal fans of generic polyhedra. Points in the interior of $\mathop{\mathrm{cone}}\Gamma$ correspond to the normal fans of full-dimensional polyhedra (see Proposition 8.2), and points on the boundary of $\mathop{\mathrm{cone}}\Gamma$ correspond to generalised fans. We illustrate this next.
@@
Figure 1.
@@
In general, the topology of real moment-angle manifolds $\mathcal{R}_{\mathcal{K}}$ and, in particular, intersections of quadrics of the form $\gamma_{1}x_{1}^{2}+\cdots+\gamma_{m}x_{m}^{2}=\delta$ is quite complicated. See [13, Chapters 4 and 6].
@@ 9. Convex hulls and links of special quadrics
We therefore refer to $\Gamma^{\prime}=\{\gamma^{\prime}_{1},\ldots,\gamma^{\prime}_{m}\}$ as a point configration in the affine space $(V^{\prime})^{*}$, as opposed to the vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$. Given a subset $I\subset[m]$, we denote by $\mathop{\mathrm{conv}}\Gamma^{\prime}_{I}$ the convex hull of the points $\gamma^{\prime}_{i}$, $i\in I$, in the affine space $(V^{\prime})^{*}$.
@@
Condition (2) in Theorem 9.2 is precisely the imbrication condition of Bosio, see [9, Lemma 1.1, Definition 1.8]. As a further specification, we obtain the following version of Theorem 8.7:
@@
Conditions (a) and (b) in Theorem 9.3 are known as the Siegel condition and the weak hyperbolicity condition, see e. g. [27, §I]. The manifold given by (9.3) is called the link of a system of special real quadrics [10, Definition 0.1]. It is diffeomorphic to the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$, where $\mathcal{K}$ is dual to the simple (generic) polytope $P$ given by
@@
Figure 2.
@@
The simplicial complex $\mathcal{K}$ (9.2) here is the boundary of pentagon with $1$-simplices $\{1,2\}$, $\{2,3\}$, $\{3,4\}$, $\{4,5\}$ and $\{5,1\}$. The intersection of quadrics (9.4) is homeomorphic to the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$, which is an oriented surface of genus $5$ (see [13, Proposition 4.1.8]). The associated polytope is given by (8.1), where $\sum_{i=1}^{5}b_{i}\gamma_{i}=\begin{pmatrix}\mathbf{0}\\ 1\end{pmatrix}$. One choice is $b_{1}=b_{3}=b_{4}=b_{5}=\frac{1}{4}$ and $b_{2}=0$. It is easy to see that $P$ is a pentagon.
@@
10. Holomorphic exponential actions and moment-angle manifolds
@@
A vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$ can be viewed as a vector configuration in the complex space $\widetilde{V}^{*}$. We continue to assume that he real span of $\Gamma$ is the whole of $V^{*}$.
@@
The stabilizer $\widetilde{V}_{{\textbf{{z}}}}$ of a point ${\textbf{{z}}}=(z_{1},\ldots,z_{m})\in\mathbb{C}^{m}$ under the action (10.1) is zero if $\mathbb{R}\langle\gamma_{i}\colon z_{i}\neq 0\rangle=V^{*}$.
@@
Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ and $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ be a pair of Gale dual vector configurations in $V^{*}$ and $W^{*}$, respectively, and let $\mathcal{K}$ be a simplicial complex on $[m]$. The holomorphic exponential action $\widetilde{V}\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, in which case the action is also free.
@@
The fact that the action $\widetilde{V}\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ defined by a fan data $\{\mathcal{K},\mathrm{A}\}$ is free follows from Proposition 10.2. ∎
@@
If $\mathbb{R}\langle\gamma_{1},\gamma_{2}\rangle=\mathbb{R}^{2}$, then the restriction of the action to (×$(\mathbb{C}^{\times})^{2}$ is free and proper by Theorem 10.3 (note that (×$(\mathbb{C}^{\times})^{2}=U_{\mathbb{C}}(\mathcal{K})$ for $\mathcal{K}=\{\varnothing\}$ on the two-element set). The quotient is given by
@@
If $\gamma_{1}=\alpha\gamma_{2}$ with $\alpha\in\mathbb{R}\setminus\mathbb{Q}$, then the action $\widetilde{V}\times\mathbb{C}^{2}\to\mathbb{C}^{2}$ is still free, so the converse of Proposition 10.1 does not hold. However, this action is not proper, as the orbit of the point $(1,1)\in(\mathbb{C}^{\times})^{2}$ is not closed. The quotient (×$(\mathbb{C}^{\times})^{2}/\widetilde{V}$ is non-Hausdorff.
Finally, if $\gamma_{1}=\alpha\gamma_{2}$ with $\alpha\in\mathbb{Q}$, then the action $\widetilde{V}\times\mathbb{C}^{2}\to\mathbb{C}^{2}$ is not free, as any $v\in\widetilde{V}$ such that $\gamma_{1}v=2\pi ik_{1}$ and $\gamma_{2}v=2\pi ik_{2}$ with integer $k_{1},k_{2}$ acts trivially. If at least one of $\gamma_{1},\gamma_{2}$ is nonzero, the quotient (×$(\mathbb{C}^{\times})^{2}/\widetilde{V}$ is $\mathbb{C}^{\times}$.
@@
It is a CW-subcomplex in the product of $m$ discs $(D^{2})^{m}$ with the induced action of $m$-torus $T^{m}=(S^{1})^{m}$. When $\mathcal{K}$ is a simplicial subdivision of an $(n-1)$-sphere, i. e. $|\mathcal{K}|\cong S^{n-1}$, the moment-angle complex $\mathcal{Z}_{\mathcal{K}}$ is a closed topological $(m+n)$-dimensional manifold, called the moment-angle manifold [13, Theorem 4.1.4]. The topology of moment-angle manifolds can be remarkably complicated, see [13, Chapter 4].
If $\{\mathcal{K},\mathrm{A}\}$ is a fan data, then the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ can be identified with the quotient $U_{\mathbb{C}}(\mathcal{K})/V$ by the real exponential action
@@
This is proved in exactly the same way as Theorem 7.3. (Also note that the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ is homeomorphic to the real moment-angle manifold $\mathcal{R}_{D(\mathcal{K})}$, where $D(\mathcal{K})$ is the double of $\mathcal{K}$, see [4] or [13, Corollary 4.8.14].)
@@
Assume the data $\{\mathcal{K},\mathrm{A}\}$ defines a complete simplicial fan $\Sigma$, and let $\widetilde{V}\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ be the corresponding holomorphic exponential action. Then there is a homeomorphism
@@
We have an inclusion of pairs $(D^{2},S^{1})\hookrightarrow(\mathbb{C},\mathbb{C}^{\times})$, which induces an inclusion $\mathcal{Z}_{\mathcal{K}}=(D^{2},S^{1})^{\mathcal{K}}\hookrightarrow(\mathbb{C},\mathbb{C}^{\times})^{\mathcal{K}}=U_{\mathbb{C}}(\mathcal{K})$. So the required homemorphism will follow from the fact that the $\widetilde{V}$-orbit of any point ${\textbf{{z}}}=(z_{1},\ldots,z_{m})\in U_{\mathbb{C}}(\mathcal{K})$ intersects $\mathcal{Z}_{\mathcal{K}}$ at a single point.
@@
Clearly, if the data $\{\mathcal{K},\mathrm{A}\}$ defines an $n$-dimensional fan $\Sigma$, then $\{\mathcal{K}_{+},\mathrm{A}_{+}\}$ defines the same fan. In this case the action above restricted to $U_{\mathbb{C}}(\mathcal{K}_{+})=U_{\mathbb{C}}(\mathcal{K})\times\mathbb{C}^{\times}$ is free and proper, and the quotient $U_{\mathbb{C}}(\mathcal{K}_{+})/V_{+}$ is homeomorphic to $\mathcal{Z}_{\mathcal{K}_{+}}=\mathcal{Z}_{\mathcal{K}}\times S^{1}$.
@@
When $\Sigma$ is the normal fan of a polytope, the quotient $U_{\mathbb{C}}(\mathcal{K})/\widetilde{V}\cong\mathcal{Z}_{\mathcal{K}}$ can be realised as a nondegenerate intersection of Hermitian quadrics.

### v2

@@ 8.3. From quadrics to polyhedra
Furthermore, the data $\{\mathcal{K},\mathrm{A}\}$ define the normal fan of a generic polyhedron $P$ given by (8.1), where $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ is the Gale dual configuration in $W^{*}$ and the numbers $b_{i}$ are defined by the expression $\delta=\sum_{i=1}^{m}b_{i}\gamma_{i}$.
@@
By definition, $\mathcal{K}$ is a pure simplicial complex of dimension $m-k-1$. Condition (b) implies that if $\delta\in\mathop{\mathrm{cone}}\Gamma_{J}$, then $\delta\in\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{J}$, $\mathop{\mathrm{cone}}\Gamma_{J}$ is a full-dimensional cone (of dimension $k$) and its relative interior is the interior. Indeed, suppose $\delta\notin\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{J}$, then $\delta$ lies in a face of $\mathop{\mathrm{cone}}\Gamma_{J}$, which is a cone of dimension $<k$. By the Carathéodory Theorem [45, Proposition 1.15], this cone is generated by subset of $\Gamma$ of cardinality $<k$, in contradiction with (b).
@@
The decomposition of $\mathop{\mathrm{cone}}\Gamma$ into cones (or chambers) $C(\delta)$ is called the GKZ decomposition (after Gelfand, Kapranov and Zelevinsky), and $\Sigma_{\Gamma}$ is called the GKZ fan, or the secondary fan of $\Gamma$, see [22, Chapter 7] and also [36], [1, §2.2.2]. It encodes generalised normal fans with cone generators chosen among the Gale dual configuration $\mathrm{A}$. The relative interiors of maximal cones of $\Sigma_{\Gamma}$ correspond to the normal fans of generic polyhedra. Points in the interior of $\mathop{\mathrm{cone}}\Gamma$ correspond to the normal fans of full-dimensional polyhedra (see Proposition 8.2), and points on the boundary of $\mathop{\mathrm{cone}}\Gamma$ correspond to generalised fans. We illustrate this next.
@@
Figure 1. GKZ decomposition and the corresponding normal fans
@@
In general, the topology of real moment-angle manifolds $\mathcal{R}_{\mathcal{K}}$ and, in particular, intersections of quadrics of the form $\gamma_{1}x_{1}^{2}+\cdots+\gamma_{m}x_{m}^{2}=\delta$ is quite complicated. See [14, Chapters 4 and 6].
@@ 9. Convex hulls and links of special quadrics
We therefore refer to $\Gamma^{\prime}=\{\gamma^{\prime}_{1},\ldots,\gamma^{\prime}_{m}\}$ as a point configuration in the affine space $(V^{\prime})^{*}$, as opposed to the vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$. Given a subset $I\subset[m]$, we denote by $\mathop{\mathrm{conv}}\Gamma^{\prime}_{I}$ the convex hull of the points $\gamma^{\prime}_{i}$, $i\in I$, in the affine space $(V^{\prime})^{*}$.
@@
Condition (2) in Theorem 9.2 is precisely the imbrication condition of Bosio, see [10, Lemma 1.1, Definition 1.8]. As a further specification, we obtain the following version of Theorem 8.7:
@@
Conditions (a) and (b) in Theorem 9.3 are known as the Siegel condition and the weak hyperbolicity condition, see e. g. [34, §I]. The manifold given by (9.3) is called the link of a system of special real quadrics [11, Definition 0.1]. It is diffeomorphic to the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$, where $\mathcal{K}$ is dual to the simple (generic) polytope $P$ given by
@@
Figure 2. Five-vector configuration and its Gale dual
@@
The simplicial complex $\mathcal{K}$ (9.2) here is the boundary of pentagon with $1$-simplices $\{1,2\}$, $\{2,3\}$, $\{3,4\}$, $\{4,5\}$ and $\{5,1\}$. The intersection of quadrics (9.4) is homeomorphic to the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$, which is an oriented surface of genus $5$ (see [14, Proposition 4.1.8]). The associated polytope is given by (8.1), where $\sum_{i=1}^{5}b_{i}\gamma_{i}=\begin{pmatrix}\mathbf{0}\\ 1\end{pmatrix}$. One choice is $b_{1}=b_{3}=b_{4}=b_{5}=\frac{1}{4}$ and $b_{2}=0$. It is easy to see that $P$ is a pentagon.
@@
10. Holomorphic exponential actions and complex-analytic structures on moment-angle manifolds
@@
A vector configuration $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$ can be viewed as a vector configuration in the complex space $\widetilde{V}^{*}$. We continue to assume that the real span of $\Gamma$ is the whole of $V^{*}$.
@@
The stabiliser $\widetilde{V}_{{\textbf{{z}}}}$ of a point ${\textbf{{z}}}=(z_{1},\ldots,z_{m})\in\mathbb{C}^{m}$ under the action (10.1) is zero if $\mathbb{R}\langle\gamma_{i}\colon z_{i}\neq 0\rangle=V^{*}$.
@@
Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ and $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ be a pair of Gale dual vector configurations in $V^{*}$ and $W^{*}$, respectively, and let $\mathcal{K}$ be a simplicial complex on $[m]$. The holomorphic exponential action $\widetilde{V}\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, in which case the action is also free.
@@
The fact that the action $\widetilde{V}\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ defined by a triangulated configuration $\{\mathcal{K},\mathrm{A}\}$ is free follows from Proposition 10.2. ∎
@@
Let $\mathcal{K}=\{\varnothing\}$ on the two-element set $[2]=\{1,2\}$, so that $U_{\mathbb{C}}(\mathcal{K})=(\mathbb{C}^{\times})^{2}$. If $\mathbb{R}\langle\gamma_{1},\gamma_{2}\rangle=\mathbb{R}^{2}$, then the restriction of the action to (×$(\mathbb{C}^{\times})^{2}$ is free and proper by Theorem 10.3 . The quotient is given by
@@
If $\gamma_{1}=\alpha\gamma_{2}$ with $\alpha\in\mathbb{R}\setminus\mathbb{Q}$, then the action $\widetilde{V}\times(\mathbb{C}^{\times})^{2}\to(\mathbb{C}^{\times})^{2}$ is still free, so the converse of Proposition 10.1 does not hold. However, this action is not proper, as the orbit of the point $(1,1)\in(\mathbb{C}^{\times})^{2}$ is not closed. The quotient (×$(\mathbb{C}^{\times})^{2}/\widetilde{V}$ is non-Hausdorff.
Finally, if $\gamma_{1}=\alpha\gamma_{2}$ with $\alpha\in\mathbb{Q}$, then the action $\widetilde{V}\times(\mathbb{C}^{\times})^{2}\to(\mathbb{C}^{\times})^{2}$ is not free, as any $v\in\widetilde{V}$ such that $\gamma_{1}v=2\pi ik_{1}$ and $\gamma_{2}v=2\pi ik_{2}$ with integer $k_{1},k_{2}$ acts trivially. If at least one of $\gamma_{1},\gamma_{2}$ is nonzero, the quotient (×$(\mathbb{C}^{\times})^{2}/\widetilde{V}$ is $\mathbb{C}^{\times}$.
@@
It is a CW-subcomplex in the product of $m$ discs $(D^{2})^{m}$ with the induced action of $m$-torus $T^{m}=(S^{1})^{m}$. When $\mathcal{K}$ is a simplicial subdivision of an $(n-1)$-sphere, i. e. $|\mathcal{K}|\cong S^{n-1}$, the moment-angle complex $\mathcal{Z}_{\mathcal{K}}$ is a closed topological $(m+n)$-dimensional manifold, called the moment-angle manifold [14, Theorem 4.1.4]. The topology of moment-angle manifolds can be remarkably complicated, see [14, Chapter 4].
If $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, then the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ can be identified with the quotient $U_{\mathbb{C}}(\mathcal{K})/V$ by the real exponential action
@@
This is proved in exactly the same way as Theorem 7.3. (Note also that the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}$ is homeomorphic to the real moment-angle manifold $\mathcal{R}_{D(\mathcal{K})}$, where $D(\mathcal{K})$ is the double of $\mathcal{K}$, see [4] or [14, Corollary 4.8.14].)
@@
Assume that $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining a complete simplicial fan $\Sigma$, and let $\widetilde{V}\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ be the corresponding holomorphic exponential action. Then there is a homeomorphism
@@
We have an inclusion of pairs $(D^{2},S^{1})\hookrightarrow(\mathbb{C},\mathbb{C}^{\times})$, which induces an inclusion $\mathcal{Z}_{\mathcal{K}}=(D^{2},S^{1})^{\mathcal{K}}\hookrightarrow(\mathbb{C},\mathbb{C}^{\times})^{\mathcal{K}}=U_{\mathbb{C}}(\mathcal{K})$. So the required homeomorphism will follow from the fact that the $\widetilde{V}$-orbit of any point ${\textbf{{z}}}=(z_{1},\ldots,z_{m})\in U_{\mathbb{C}}(\mathcal{K})$ intersects $\mathcal{Z}_{\mathcal{K}}$ at a single point.
@@
Clearly, if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining an $n$-dimensional fan $\Sigma$, then $\{\mathcal{K}_{+},\mathrm{A}_{+}\}$ defines the same fan. In this case, the exponential action of $V_{+}$ above restricted to $U_{\mathbb{C}}(\mathcal{K}_{+})=U_{\mathbb{C}}(\mathcal{K})\times\mathbb{C}^{\times}$ is free and proper. When the fan is complete, the quotient $U_{\mathbb{C}}(\mathcal{K}_{+})/V_{+}$ is homeomorphic to $\mathcal{Z}_{\mathcal{K}_{+}}=\mathcal{Z}_{\mathcal{K}}\times S^{1}$.
@@
Proposition 10.10.
A complex moment-angle manifold $\mathcal{Z}_{K}$ is non-Kähler, unless it is a holomorphic torus.
Proof.
Suppose that $\mathcal{Z}_{K}$ is Kähler of complex dimension $d$. Then the cohomology class $[\omega]\in H^{2}(\mathcal{Z}_{K})$ of the Kähler form satisfies $[\omega]^{d}\neq 0$, as $\mathcal{Z}_{K}$ is compact.
If $\mathcal{K}$ does not have ghost vertices, then it is easy to see using (10.3) and the appropriate cell decomposition that $\mathcal{Z}_{K}$ is 2-connected (see [14, Proposition 4.3.5]). In general, $\mathcal{Z}_{K}$ is diffeomorphic to a product of a 2-connected manifold and a torus (see Construction 10.8). Hence, it can have a 2-dimensional cohomology class satisfying $[\omega]^{d}\neq 0$ only if the 2-connected factor is a point, so $\mathcal{K}$ consists entirely of ghost vertices and $\mathcal{Z}_{K}$ is a $d$-dimensional holomorphic torus. ∎
When $\Sigma$ is the normal fan of a polytope, the quotient $U_{\mathbb{C}}(\mathcal{K})/\widetilde{V}\cong\mathcal{Z}_{\mathcal{K}}$ can be realised as a nondegenerate intersection of Hermitian quadrics: